### v1

# Interpretable Embeddings of Speech Enhance and Explain Brain Encoding Performance of Audio Models

Abstract
Self-supervised speech models (SSMs) are increasingly hailed as more powerful computational models of human speech perception than models based on traditional hand-crafted features. However, since their representations are inherently black-box, it remains unclear what drives their alignment with brain responses. To remedy this, we built linear encoding models from six interpretable feature families: mel-spectrogram, Gabor filter bank features, speech presence, phonetic, syntactic, and semantic Question-Answering features, and contextualized embeddings from three state-of-the-art SSMs (Whisper, HuBERT, WavLM), quantifying the shared and unique neural variance captured by each feature class. Contrary to prevailing assumptions, our interpretable model predicted electrocorticography (ECoG) responses to speech more accurately than any SSM. Moreover, augmenting SSM representations with interpretable features yielded the best overall neural predictions, significantly outperforming either class alone.
Further variance-partitioning analyses revealed previously unresolved components of SSM representations that contribute to their neural alignment: 1. Despite the common assumption that later layers of SSMs discard low-level acoustic information, these models compress and preferentially retain frequency bands critical for neural encoding of speech (100–1000 Hz). 2. Contrary to previous claims, SSMs encode brain-relevant semantic information that cannot be reduced to lower-level features, improving with context length and model size. These results highlight the importance of using refined, interpretable features in understanding speech perception.
1 Introduction
Self-supervised models have emerged as increasingly popular tools for studying how the brain processes language and speech. Numerous studies have employed large language models (LLMs) and demonstrated their effectiveness in predicting neural responses to narrative stimuli by linearly probing their internal representations (Toneva and Wehbe 2019; Goldstein et al. 2022; Caucheteux and King 2022; Antonello, Vaidya, and Huth 2023; Hosseini et al. 2024; Mischler et al. 2024). Similarly, self-supervised speech models (SSMs) have been used to investigate neural representations underlying speech perception. These studies commonly conclude that, like LLMs, SSMs are highly effective in predicting brain responses to speech, often outperforming their interpretable baselines (Vaidya, Jain, and Huth 2022; Anderson, Davis, and Lalor 2024; Goldstein et al. 2025; Tuckute et al. 2023). Consequently, these findings are interpreted as evidence that SSMs are more powerful and brain-like models of speech perception, with their superiority often attributed to their ability to integrate contextual information over time (Anderson, Davis, and Lalor 2024; Goldstein et al. 2025).
However, despite the apparent success of SSMs in predicting brain responses, several critical challenges remain in drawing meaningful neuroscientific conclusions from these results. First, the goal of cognitive neuroscience is not to build computational models with the best predictive power, but to understand the mechanisms underlying such neural computation. While there has been growing effort to investigate what these models encode, much of the existing work has focused on characterizing their internal representations (Pasad, Chou, and Livescu 2022; Pasad, Shi, and Livescu 2023; Pasad et al. 2024; Ashihara et al. 2023; Shen et al. 2023; Choi and Yeo 2022; Choi et al. 2024), an important but distinct goal from identifying which aspects of their representations are relevant for neural processing. Techniques such as Canonical Correlation Analysis (CCA), Representational Similarity Analysis (RSA), and linear probing with interpretable embeddings are commonly used to probe the internal structure of SSMs. However, these methods typically assess aggregate similarity across representational spaces and do not isolate specific components that are mechanistically meaningful for brain function. As a result, even when combined with strong neural prediction performance, such findings offer limited guidance for understanding how model representations map onto neural processes and may be prone to misinterpretation. Second, the predictive performance of SSMs may potentially be explained by missed, yet simple confounds. In the context of LLMs, Hadidi et al. showed that most of the neural variance explained by GPT2-XL can be explained by simple confounds such as word-rate and positional information (Hadidi et al. 2025). This raises the concern that SSMs, too, might be over-credited with encoding unique brain-relevant contextual information when, in fact, their performance is driven by simpler, interpretable confounds. Collectively, these challenges underscore the importance of disentangling what properties drive the predictive performance of SSMs and assessing the extent to which their contextual representations capture unique neural variance beyond a broad set of interpretable features.
A detailed examination of the existing empirical evidence further highlights these concerns. While there is robust evidence that SSMs yield better alignment in higher-order brain regions than simple acoustic baselines (Vaidya, Jain, and Huth 2022; Li et al. 2023), the nature of the representations driving this alignment remains unclear, particularly with regard to whether they encode brain-relevant semantic information. For example, Anderson et al. (Anderson, Davis, and Lalor 2024) found that combining representations from Whisper’s later layers with those of GPT-2 did not improve alignment with EEG data compared to Whisper alone, suggesting that Whisper already encodes lexical content relevant for neural processing. In contrast, Oota et al. (Oota et al. 2024) found that after regressing out low-level acoustic and speech features, the representations of the SSMs showed no significant alignment with high-level language regions of the brain, indicating that these models may lack higher-order semantic information. We argue that these contrasting findings may stem from the indirect nature of current methodologies, which often involve inferring properties of one black box (SSMs) using another (the brain) without the use of interpretable controls, making interpretation inherently challenging.
In contrast to previous work, this study directly addresses these open questions using a variance partitioning approach applied to hand-crafted, interpretable features spanning low-level acoustic to high-level lexical domains. Specifically, we show that: 1. Handcrafted interpretable features can outperform SSM representations in predicting brain responses. Combining interpretable features and SSM representations leads to the best performance. 2. SSM representations retain brain-relevant low-level acoustic information across layers, with acoustic features becoming increasingly compact to accommodate higher-level representations. 3. SSMs encode semantic information that is relevant to neural processing, which improves with context length and model size.
Figure 1: Summary of the methodologies. (a) This study constructs six sets of interpretable features—mel-spectrograms, Gabor filter bank features, speech presence, phonetic, syntactic, and semantic features—and compares them to representations from three state-of-the-art self-supervised speech models (SSMs): Whisper, HuBERT, and WavLM. The comparison is based on each representation’s ability to predict neural responses, measured through Pearson’s $r$ and variance explained $R^{2}$. (b) An example of each linguistic feature set. The word is represented as phonetic (ARPAbet), syntactic (Part-of-Speech Tag, Dependency Parsing, stop word or not), and semantic (Question Answering) features. (c) Diagram of the variance partitioning approach.
2 Methods
Fig. 1 presents a summary of the methodologies employed in this research. All results can be reproduced using the code available at: https://github.com/rikishimizu/interpretable˙speech
2.1 Datasets
All results were derived using open-source data contained in the Podcast ECoG Dataset (Zada et al. 2025). It contains electrocorticography (ECoG) recordings from nine patients with epilepsy who listened to a 30-minute excerpt of the podcast This American Life (“So a Monkey and a Horse Walk Into a Bar: Act One—Monkey in the Middle”). Of the 1,330 intracranial electrodes originally implanted, 1,268 were retained after excluding contacts with imprecise anatomical localization or excessive noise in their power-spectral density. Electrode positions were projected onto a template cortical surface for visualization. Full pre-processing details are provided in (Zada et al. 2025).
We focused on broadband high-gamma activity (70–200 Hz), extracted with a fourth-order zero-phase Butterworth band-pass IIR filter and converted to analytic amplitude. The resulting envelopes were down-sampled to 4 Hz and z-scored.
2.2 Interpretable auditory features
We derived six families of interpretable predictors: (i) mel-spectrograms, (ii) Gabor filter bank (GBFB) features, (iii) speech presence, (iv) phonetic features, (v) semantic features, and (vi) syntactic features (Fig. 1a). These features reflect the language processing hierarchy in speech comprehension (Heer et al. 2017; Keshishian et al. 2023). For phonetic, semantic, and syntactic features, the feature vector at time $t$ s was computed by summing up the embeddings for all words or phonemes whose offset time lies in [$t-0.5$, $t$]s. Full details of each feature processing are available in Section S2.
Mel-spectrogram
We computed mel-spectrograms from 0–8 kHz with 32 mel bins and further converted them to decibels using librosa (McFee et al. 2015).
Gabor bank filter features (GBFB)
Spectro-temporal modulation features were extracted with the Gabor filter-bank implementation of Schädler et al (Schädler, Meyer, and Kollmeier 2012).
Speech presence
Speech probability was estimated with EfficientAT, a CNN-based audio tagger (Schmid, Koutini, and Widmer 2023).
Phonetic features
Each word in the transcript was converted to its ARPAbet phoneme representations using nltk.corpus.cmudict (Bird and Loper 2004). Every phoneme was one-hot encoded over the 39 ARPAbet symbols.
Semantic features
Following Benara et al. 2024, we obtained 168-dimensional question-answering (QA) semantic embeddings by prompting GPT-4 (gpt-4-0125-preview; OpenAI 2023) with 56 yes/no questions about each word’s meaning (e.g., “Does the word/sentence involve an expression of personal values or beliefs?”) with 3 different context lengths (word-level, 1.5 s, and 3 s). For the 1.5/3-second context length, the model was additionally provided with the words spoken in the 1.5/3 seconds preceding each target word when answering the questions. Because of how this feature is constructed, its effective context window can extend up to approximately 4.75 seconds—accounting for the 3-second input context, the additional 0.5-second summing window, and the maximum word duration (about 1.25 seconds for words like “approximately”). A full list of questions is provided in Table S2.
Syntactic features
Following Zada et al. 2025, each word was first converted into 96-dimensional embeddings, comprising 50 part-of-speech tags, 45 dependency labels, and a stop/non-stop label using the spaCy (Honnibal et al. 2020).
2.3 Speech Audio Models
We used three speech models: Whisper (Radford et al. 2022), HuBERT (Hsu et al. 2021), and WavLM (Chen et al. 2022). (Fig. 1a)
For this work, unless otherwise specified, we used the largest variants of these models: Whisper Large V1 (638M Parameter, 32 layers), HuBERT X-large (964M Parameter, 48 layers), and WavLM X-large (317M Parameter, 24 layers). Table S1 shows the summary of the architecture and training objectives for each model. To generate the embeddings with a context length of $Ctx$ seconds, we applied a sliding window of $Ctx$ seconds and a stride of 250 ms to the input audio. For each window, the representation at a given layer was defined as the hidden state of the final token. For Whisper, we only used hidden states from the encoder. In this study, $Ctx$ was varied across the following values: 0.5, 1, 2.5, 5, 10, 20, and 30 seconds. PCA with $95\%$ variance was applied to each model’s embeddings to reduce dimensionality.
2.4 Encoding Model Construction
After processing, the high gamma signals we predict have shape $[T,N_{ch}]$ where $T$ is the number of time points (i.e., 7200) and $N_{ch}$ is the number of channels (i.e., 1268). To estimate the model performance, we used a 5-fold cross-validation scheme, splitting the signals into 5 temporally contiguous chunks. For each fold, the model was trained on $80\%$ of the data and tested on the remaining $20\%$. Within the training folds, we employed a 4-fold nested cross-validation scheme to optimize the model parameters. Unless otherwise noted—specifically in analyses examining the impact of context size on encoding performance—the optimal context size was selected and used for each speech audio model.
Banded ridge regression
To create a joint encoding model with features from different feature spaces and dimensionalities, we employed banded ridge regression, which has been shown to improve neural encoding performance when using diverse feature spaces (Nunez-Elizalde, Huth, and Gallant 2019). Unlike ridge regression with spherical Gaussian priors, banded ridge regression optimizes the regularization parameter alpha separately for each feature space. We used the himalaya library by Dupre La Tour et al. (Dupré la Tour et al. 2022). Each feature was z-scored before applying the regression model. Unless otherwise specified, we varied alpha over 20 logarithmically spaced points between $10^{0}$ and $10^{3}$.
Performance Metrics
To measure the performance of the encoding models, we used correlation (Pearson’s $r$) and the variance explained $R^{2}$, two standard metrics for evaluating encoding models (Vaidya, Jain, and Huth 2022; Anderson, Davis, and Lalor 2024; Goldstein et al. 2025; Tuckute et al. 2023; Li et al. 2023). Generally, we used Pearson’s $r$ to compare the encoding model performance of different representations, while the variance explained $R^{2}$ was used to quantify the unique and shared contributions of different features for neural prediction. Negative $R^{2}$ values were clipped to 0, since negative explained variance lacks interpretability.
Variance Partitioning
To quantify the shared and unique variance contributions of feature sets, we used the variance partitioning method established by de Heer et al. (Heer et al. 2017). For details, refer to Section S4 or Heer et al. 2017.
Computational Resources
All experiments and analyses were conducted on AMD EPYC 7662 CPUs and NVIDIA L40 GPUs, except for the generation of QA semantic features, which was performed using the OpenAI API. Cumulatively, the experiments required approximately 500 GPU hours.
3 Results
3.1 Interpretable models outperform and enhance SSMs in predicting brain responses
Figure 2: Interpretable model outperforms and enhances SSM representations in predicting brain responses. (a) Prediction performance of the model based on our interpretable features for all electrodes, plotted on a template brain. (b) Average Pearson’s $r$ across all electrodes for different models. For each SSM, the context size with the highest average Pearson’s $r$ was chosen. Joint Model refers to a combination of interpretable features and Whisper 32nd layer representations with 30 seconds of context. The interpretable model outperformed all SSMs, and the joint model of interpretable features and SSM embeddings achieved the highest average Pearson’s $r$. (c) Pearson’s $r$ for different models with the best layer and context size, shown for different anatomical regions (Paired t-test, FDR correction, two-sided). The interpretable model performs well across anatomical regions and significantly outperforms SSMs in lower-level regions, such as Heschl’s gyrus and the insula.
Fig. 2a shows the spatial distribution of electrode-wise correlation values for our interpretable model, mapped onto a template brain. Fig. S1 quantifies the contribution of unique variance explained $R^{2}$ by each feature set. Notably, the language processing hierarchy can be clearly observed in the anatomical distributions.
Fig. 2b shows the average Pearson’s $r$ values across all electrodes for each SSM over layers, compared with our interpretable model. For each model, the context size with the highest peak average correlation value was chosen and noted. Our interpretable model achieved an average correlation value of 0.079, outperforming the highest average correlation values of 0.073, 0.073, and 0.069 observed for WavLM, Whisper, and HuBERT, respectively ($p<0.001$, paired t-test, Fig. S2). The joint models of interpretable features with SSM representations achieved average correlation values of approximately 0.082 (0.0828 for WavLM, 0.0827 for Whisper, and 0.0821 for HuBERT), significantly outperforming models using only interpretable features ($p<0.001$, paired t-test). In agreement with the previous findings (Vaidya, Jain, and Huth 2022; Tuckute et al. 2023; Anderson, Davis, and Lalor 2024), all three models significantly outperformed the model based solely on mel-spectrogram, which achieved an average correlation of 0.057 ($p<0.001$, paired t-test). An anatomical analysis of encoding performance (Fig. 2c) shows that the interpretable model consistently outperforms all SSMs in Heschl’s gyrus and the insula—two regions known to encode low-level acoustic features (Berger et al. 2025; Kurteff et al. 2024; Khalighinejad et al. 2021; Leaver and Rauschecker 2010). Even when the comparison is applied to early layers of SSMs, this difference remains statistically significant, except for HuBERT, although the gap in the average correlation slightly narrows (Fig. S3). These results match with findings from Li et al. that concluded that HG representation is not modeled well by speech DNNs (Li et al. 2023).
Previous work showed that later layers of self-supervised models (SSMs) discard low-level acoustic details (Pasad, Chou, and Livescu 2022; Pasad et al. 2024). If the discarded details are important for predicting neural responses, combining early and late layer representations should improve predictive performance. To test this hypothesis, we evaluated joint models of representations from early and late SSM layers (Fig. 2b). Due to the high dimensionality of the combined features, we varied $\alpha$ across 40 logarithmically spaced values from $10^{0}$ to $10^{6}$. Combining Whisper layers 1 and 32 improved average Pearson’s r by 0.003 over layer 32 alone; WavLM layers 1 and 20 improved by 0.00024; and HuBERT layers 1 and 32 by 0.0025. While these gains may be diminished by overfitting due to increased dimensionality, these gains remain substantially smaller than the predictive performance of the mel-spectrogram (r = 0.056), which can be viewed as an approximate upper bound on the expected improvement if later SSM layers had entirely discarded low-level acoustic information. This suggests that later SSM layers may retain more low-level acoustic information than previously assumed, at least in forms relevant for neural processing.
3.2 SS Models compress and retain brain-relevant low-level acoustic information
Figure 3: SSMs compress and retain brain-relevant low-level acoustic information. (a) The ratio of neural variance shared by the mel-spectrogram and the SSM representations, relative to the total variance explained by the mel-spectrogram. The ratio remains consistently high across layers, suggesting that acoustic information relevant for neural processing is preserved. (b) Average Variance $R^{2}$ of the SSM representations across their feature dimensions, explained by the mel-spectrogram. This metric quantifies how much variance in SSM embeddings is allocated to encoding mel-spectrogram information, and shows a substantial decrease in later layers. (c) Variance $R^{2}$ of the mel-spectrogram [$t-0.5$, $t$]s explained by the Whisper representations over different layers. (d) Average variance explained $R^{2}$ over the 0.2 seconds [$t-0.2$, $t$]s for different frequency bins, plotted over layers. The gain/loss in percentage with respect to the 1st layer is displayed on the middle plot. On the right plot, the gain/loss percentage for the 32nd layer is shown, demonstrating the preferential preservation of speech critical frequency bands.
Inspired by the findings in the previous section, we conducted a mechanistic interpretability analysis on how low-level acoustic information is represented across the layers of the SSMs.
Fig. 3a shows the proportion of neural variance in Heschl’s gyrus and Insula shared by the mel-spectrogram and the SSM representations, relative to the total variance explained by the mel-spectrogram. This ratio reflects the extent to which each layer of an SSM encodes brain-relevant low-level acoustic information. As shown in the figure, this ratio remains consistently high across layers: for Whisper, it peaks at 90.4% and decreases to 77.5% in later layers; for WavLM, from 89.5% to 80.1%; and for HuBERT, from 96.2% to 71.9%.
To better understand how the mel-spectrogram is represented internally within each model, Fig. 3b shows the average variance explained by the mel-spectrogram when directly predicting SSM representations across layers. This metric quantifies the degree to which the internal representations of the SSMs are allocated to encode the feature. For Whisper, the average explained variance $R^{2}$ peaks at layer 1 with a value of 0.204 and drops to 0.093 in the last layer, corresponding to relative reductions of 54.4%. For WavLM, it decreases from 0.192 to 0.120 (37.5% reduction); and for HuBERT, from 0.234 to 0.055 (76.5% reduction). Notably, these reductions are substantially larger than the relatively modest declines observed in neural variance shared between the mel-spectrogram and SSM representations in Heschl’s gyrus and Insula, suggesting that while low-level acoustic information becomes increasingly compressed within the models, it remains selectively preserved in brain-relevant representations.
Next, to investigate how these models retain brain-relevant low-level acoustic information over layers, we directly predicted 0.5-second windows of mel-spectrogram from the SSM representations: that is, using the SSM representations at time t s, we predicted the entire mel-spectrogram in the [$t-0.5$, $t$]s window. Fig. 3(c) shows the variance $R^{2}$ of the mel-spectrogram explained by Whisper representations over different layers. It can be observed that in the earlier layers (0th and 8th), the Whisper representations predict all frequency bins equally well, but their predictive performance in the lowest frequency bin and higher frequency bins drops markedly in later layers. Fig. 3(d) quantifies this trend by averaging the variance explained across the [$t-0.2$, $t$]s window for each frequency bin. Frequencies in the 100-1000 Hz range exhibit a significantly smaller decrease in explained variance compared to frequencies outside this range (average: $9.9$% vs. $17.9$%). Fig. S5 shows a similar trend for WavLM and HuBERT as well. This frequency band aligns with regions of greater cortical representation ($<$ 1000 Hz) during speech perception, as previously reported in Heschl’s gyrus and surrounding areas (Moerel, Martino, and Formisano 2012; Khalighinejad et al. 2021). Together, these results suggest that while SSMs compress low-level acoustic features internally, they preferentially preserve the components most relevant to brain responses during speech perception.
3.3 SS Models Encode Brain-Relevant Semantic Information
Figure 4: SSMs encode brain-relevant semantic information. (a) Diagram of the three-way variance partitioning analysis used to assess whether SSM representations encode semantic information. (b) Unique variance explained $R^{2}$ by different feature sets in predicting GPT-2 representations, averaged across GPT-2’s feature dimensions ($p<0.001$, Wilcoxon signed-rank test). The significant unique contribution of our question-answering semantic features confirms that they encode distinct semantic information. (c) Unique neural variance explained by semantic QA features relative to non-semantic interpretable features. (d) Box plot of the values in (c), across electrodes in the inferior frontal gyrus triangularis (IFG-Tri). Median values were significantly greater than zero ($p<0.001$, Wilcoxon signed-rank test). (e) $\Delta R^{2}$, as illustrated in (a), averaged across electrodes in IFG-Tri. Left: Varying context lengths, using the largest variants of each SSM. Right: Varying model sizes, with context length fixed to 5 seconds for WavLM and HuBERT, and 30 seconds for Whisper. For each model size and context length, the layer with the highest $\Delta R^{2}$ was chosen. It shows SSMs encode brain-relevant semantic information that cannot be reduced to lower-level features, improving with context length and model size.
Fig. 4a outlines our approach for evaluating whether SSM representations encode semantic information. First, we calculate the unique neural variance $R^{2}$ explained by semantic QA features with respect to non-semantic interpretable feature sets (mel-spectrogram, GBFB, speech presence, phonetic, and syntactic features). We then assess how much this unique neural variance is reduced when SSM representations are included in the model alongside the non-semantic features. This reduction reflects the shared neural variance $R^{2}$ between semantic QA features and SSM representations that is not explained by non-semantic interpretable feature sets.
QA features encode genuinely semantic information relevant for brain responses
Before we assess SSMs, to evaluate whether our QA features actually encode semantic information, we computed the unique variance $R^{2}$ by each interpretable feature set in predicting GPT-2 representations (Fig. 4b). This step was included to ensure the quality of QA features, which depend on the accuracy of the LLM’s answers. Syntactic, semantic, and phonetic features account for the largest proportion of unique variance ($p<0.001$, Wilcoxon signed-rank test) across GPT-2’s feature dimensions. Speech probability also explains a statistically significant amount of unique variance, though its contribution is markedly smaller in comparison. In contrast, GBFB and mel-spectrogram features do not explain any significant unique variance, as expected—GPT-2 is not trained on acoustic input and thus should not encode such information. The significant unique variance explained by our semantic features supports the claim that they indeed capture meaningful semantic content.
Next, to test whether semantic information captured by QA features is relevant to neural processing, we computed the variance of the brain responses uniquely explained by semantic features relative to non-semantic interpretable features (Fig. 4c). Semantic features had the largest unique variance explained in the inferior frontal gyrus pars triangularis (IFG-Tri) - a part of Broca’s area that was previously found to be important for processing the semantic meaning of words (Newman et al. 2003; Vigneau et al. 2006; Friederici, Opitz, and von Cramon 2000). As shown in Fig. 4(d), the median of the unique neural variance explained by semantics in IFG-Tri was significantly above 0 ($p<0.001$, Wilcoxon signed-rank test). These results confirm that our semantic QA features are not only semantically meaningful but also relevant to brain regions involved in semantic comprehension.
SS models encode semantic information that is relevant for brain responses
The left plot of Fig. 4e shows $\Delta R^{2}$, the shared neural variance between semantic QA features and SSM representations that is not accounted for by non-semantic interpretable features, across different context sizes. Averaged across all electrodes in IFG-Tri, $\Delta R^{2}$ was significantly above 0 ($p<0.001$, Wilcoxon signed-rank test). For WavLM and HuBERT, it increases up to a 5-second context window, consistent with the 4.75-second intrinsic context of our semantic QA features. For Whisper, the increase extended up to 30 seconds, potentially because Whisper was trained strictly with 30 seconds of audio. The right plot in Fig. 4e shows the same metric, but for varying model sizes with fixed context size. Here, the average $\Delta R^{2}$ increases with model size, suggesting that larger models encode more brain-relevant semantic information.
4 Discussion
Recent papers (Anderson, Davis, and Lalor 2024; Goldstein et al. 2025) have claimed a paradigm shift from context-invariant handcrafted features to SSM embeddings in building computational models of speech perception. However, our findings caution against treating SSM embeddings as a standalone complete model for speech perception. We showed that interpretable models alone can outperform SSMs in predicting brain responses, especially in brain regions responsible for processing low-level acoustic information. Our results showed that the combination of SSM embeddings and interpretable features achieved the best predictive performance, highlighting the importance of the continued use of interpretable features in building computational models of human speech perception.
One of our main contributions is to explicitly demonstrate that SSMs preferentially preserve low-level acoustic information within the speech critical bands (100 – 1000 Hz). Previous studies (Pasad, Chou, and Livescu 2022; Pasad, Shi, and Livescu 2023) on the layer-wise representation in SSMs analyzed acoustic information more broadly, without examining frequency-specific selectivity. To the best of our knowledge, the closest related findings are: Wang et al. (Wang and Yamagishi 2022) showed that speech anti-spoofing systems based on SSMs were relatively robust to band-stop filtering of high frequencies but not to filtering in the 0-0.8 kHz range. Our results build on these observations by providing direct evidence of frequency-selective representations within SSMs, showing that the preservation of low-frequency components is not only critical for performance but also intrinsically encoded in the model’s internal representations.
Whether SSMs encode brain-relevant semantics is debated, with evidence supporting (Pasad, Chou, and Livescu 2022; Chen et al. 2024; Anderson, Davis, and Lalor 2024) and challenging (Choi et al. 2024; Oota et al. 2024). Our method employs an interpretable semantic embedding to directly test for the shared neural variance with SSMs, while controlling for lower-level acoustic and linguistic information. We show that SSM representations significantly reduce the unique variance explained by the semantic embedding, providing strong evidence that they indeed encode brain-relevant semantic information. However, an important nuance remains: SSMs account for only about 40% of the unique variance explained by the semantic features, indicating that their semantic representations remain partial. Recent studies (Moussa, Klakow, and Toneva 2025; Vattikonda et al. 2025) have shown that fine-tuning SSMs on fMRI data can improve their alignment with language regions of the brain. Thus, while our findings affirm the presence of semantic information in SSM representations, they also suggest room for improvement.
Limitations
1. Our findings are based on data from nine clinical ECoG participants who listened to a single 30‑min English podcast. While the high temporal resolution of ECoG and broad electrode coverage across brain regions make this dataset particularly well-suited for the present study, replication using larger and more diverse datasets will be important for assessing the generalizability of these results. 2. While we controlled for several lower-level features in our analysis of semantic encoding in SSMs, we cannot entirely rule out the possibility that unaccounted confounds contribute to the shared variance between SSM representations and our semantic features. However, the observed effect of context size provides some evidence against this interpretation, as lower-level features typically operate over shorter integration windows and are therefore less sensitive to extended context. 3. While our interpretable features account for most of the neural variance captured by SSM representations, a statistically significant portion remains unexplained. Identifying the nature of this residual variance is an important direction for future research.
Acknowledgements
The authors thank the members of the lab for reviewing our manuscript and offering helpful feedback. We also thank Zada et al. for making a valuable ECoG dataset public.
Supplementary Materials for ”Interpretable Embeddings of Speech Enhance and Explain Brain Encoding Performance of Audio Models”
S1 SSM Feature Extraction
Representations from self-supervised speech models (SSMs) were extracted using the code provided by Antonello et al. (available at https://github.com/HuthLab/encoding-model-scaling-laws). The input audio file, totaling 1800 seconds, was divided into five contiguous 360-second segments before being fed into the SSMs. This segmentation was performed to prevent data leakage between training and test folds in the 5-fold cross-validation scheme.
We used the largest variants of each SSM (Whisper, WavLM, and HuBERT) because prior findings from Antonello et al. (Antonello, Vaidya, and Huth 2023) showed that the encoding performance of the brain by SSMs increased logarithmically with the number of model parameters. The details of each model are shown in Table.S1.
S2 Interpretable auditory features
Full details of the processing procedures for each interpretable feature set are provided below.
Mel-spectrogram
We computed mel-spectrograms from 0–8 kHz with 32 mel bins and further converted them to decibels using librosa (McFee et al. 2015). To create a feature vector for time $t$ s, we concatenated the frames spanning [$t-0.5$, $t$]s, yielding a 672-dimensional vector. Dimensionality was reduced by principal-component analysis (PCA), retaining 95% of the variance.
Gabor bank filter features (GBFB)
Spectro-temporal modulation features were extracted with the Gabor filter-bank implementation of Schädler et al (Schädler, Meyer, and Kollmeier 2012). The 455-dimensional features, computed at 100 Hz, were down-sampled to 4 Hz to align with the neural data; PCA (95% variance) was applied prior to modelling.
Speech presence
Speech probability was estimated with EfficientAT, a CNN-based audio tagger (Schmid, Koutini, and Widmer 2023). To produce a feature scalar for time $t$ s, we fed the model the audio spanning [$t-0.5$, $t$]s.
Phonetic features
Each word in the transcript was converted to its ARPAbet phoneme representations using nltk.corpus.cmudict (Bird and Loper 2004). Word intervals were evenly subdivided among their $N$ phones, assigning on- and offset times to every phoneme. Every phoneme was one-hot encoded over the 39 ARPAbet symbols. To create a feature vector for time $t$ s, the embeddings for every phoneme whose offset time lies in [$t-0.5$, $t$]s were summed, producing a $T\times 39$ matrix.
Semantic features
Following Benara et al. 2024, we obtained 168-dimensional question-answering (QA) semantic embeddings by prompting GPT-4 (gpt-4-0125-preview; OpenAI 2023) with 56 yes/no questions about each word’s meaning (e.g., “Does the word/sentence involve an expression of personal values or beliefs?”) with 3 different context lengths (word-level, 1.5 s, and 3 s). For the 1.5/3-second context length, the model was additionally provided with the words spoken in the 1.5/3 seconds preceding each target word when answering the questions. To create a feature vector for time $t$ s, the embeddings for every word whose offset time lies in [$t-0.5$, $t$]s were summed, producing a $T\times 168$ matrix. Because of how this feature is constructed, its effective context window can extend up to approximately 4.75 seconds—accounting for the 3-second input context, the additional 0.5-second summing window, and the maximum word duration (about 1.25 seconds for words like “approximately”). A full list of questions is provided in Table S2.
Syntactic features
Following Zada et al. 2025, each word was first converted into 96-dimensional embeddings, comprising 50 part-of-speech tags, 45 dependency labels, and a stop/non-stop label using the spaCy (Honnibal et al. 2020) library and its en_core_web_lg model. To create a feature vector for time $t$ s, the embeddings for every word whose offset time lies in [$t-0.5$, $t$]s were summed, producing a $T\times 96$ matrix.
S3 Semantic Question-Answering Features
The list of all questions for generating semantic Question-Answering features is provided in Table.S2. The categories were assigned based on clustering on their cosine similarity. Questions 36-56 were not assigned any category as they were deemed less relevant for neural encoding than the first 35 questions.
Fig. S6 shows the results of category-level analysis, in which questions in QA features were divided into sub-categories (Miscellaneous, Visual, Social, Numeric, Abstract / Opinion) based on the clustering of these questions. The results show that SSMs encode semantic information broadly rather than being limited to any specific semantic sub-category, provided the information is neurally relevant.
S4 Variance Partitioning
Between two feature sets $\mathcal{A}$ and $\mathcal{B}$,
$R^{2}_{\mathcal{A}_{\textbf{unique}}}$ $=R^{2}_{\mathcal{(A,B)}}-R^{2}_{\mathcal{B}}$
$R^{2}_{\mathcal{B}_{\textbf{unique}}}$ $=R^{2}_{\mathcal{(A,B)}}-R^{2}_{\mathcal{A}}$
$R^{2}_{{}_{\textbf{shared}}}$ $=R^{2}_{\mathcal{A}}+R^{2}_{\mathcal{B}}-R^{2}_{\mathcal{(A,B)}}$
where $R^{2}_{\mathcal{A}}$, $R^{2}_{\mathcal{B}}$, and $R^{2}_{\mathcal{(A,B)}}$ denote the variance explained by models using only $\mathcal{A}$, only $\mathcal{B}$, and both $\mathcal{A}$ and $\mathcal{B}$ together, respectively. All models were fitted using banded ridge regression.
S5 Effect of context length on brain prediction performance
Among the context sizes tested, the context size with the highest average correlation was 30 seconds for Whisper and HuBERT, and 5 seconds for WavLM (Fig. S4). This variability highlights the need for caution when interpreting the influence of context length on encoding performance. For instance, Anderson et al. found that EEG encoding performance with Whisper peaked at context lengths of 5–10 seconds and interpreted this as evidence that EEG preferentially reflects the encoding of speech within that temporal window. However, it is essential to distinguish whether such patterns reflect inherent properties of neural processing or are instead artifacts of model-specific biases.
Figure S1: Unique variance explained $R^{2}$ by each feature set. The language processing hierarchy can be clearly observed in the anatomical distributions.
Figure S2: Overall Average Pearson’s $r$ across electrodes for different models with the best layer and context size. The difference was found to be statistically significant ($p<0.001$) with paired t-test.
Figure S3: Average Pearson’s $r$ in Heschl’s gyrus and Insula for the interpretable model and early/late layers of SSMs. We assessed statistical significance using paired t-tests followed by FDR correction.
Figure S4: Average Pearson’s $r$ across all electrodes for the different context sizes of SSMs. Left: Whisper, Middle: WavLM, Right: HuBERT
Figure S5: (a) Variance $R^{2}$ of mel-spectrogram [$t-0.5$, $t$]s explained by HuBERT representations over different layers. (b) Average variance explained $R^{2}$ by HuBERT over the 0.2 seconds [$t-0.2$, $t$]s for different frequency bins, plotted over layers. The gain/loss in percentage with respect to the 1st layer is displayed on the middle plot. On the right plot, the gain/loss percentage for the 37th layer is shown. (c, d): Corresponding plots for WavLM, analogous to (a) and (b), respectively.
Figure S6: Unique variance $R^{2}$ explained by 5 categories of semantic features in IFG-Tri relative to non-semantic interpretable features and the joint model with SSM representations.
Table S1: Summary of the self-supervised speech models used in this study.
Model # Layers # Dimensions # Parameters Training Objective
HuBERT X-large 48 1280 964M Masked prediction of quantised latent units from iterative k-means clustering of MFCCs
WavLM Large 24 1024 317M Same masked-unit objective as HuBERT $+$ utterance-mixing and denoising augmentation
Whisper Large 32 1280 638M Supervised multilingual encoder–decoder ASR and speech-to-text translation
Table S2: List of all questions in question-answering features and their categories. N/A means not assigned any category.
Question Category
1. Does the sentence describe a personal reflection or thought? Abstract / Opinion
2. Does the sentence contain a proper noun? Misc
3. Does the sentence describe a physical action? Visual
4. Does the sentence describe a personal or social interaction that leads to a change or revelation? Social
5. Does the sentence involve the mention of a specific object or item? Visual
6. Does the sentence involve a description of physical environment or setting? Visual
7. Does the sentence describe a relationship between people? Social
8. Does the sentence mention a specific location? Misc
9. Is time mentioned in the input? Numeric
10. Is the sentence abstract rather than concrete? Abstract / Opinion
11. Does the sentence express the narrator’s opinion or judgment about an event or character? Misc
12. Is the input related to a specific industry or profession? Misc
13. Does the sentence include dialogue? Social
14. Does the sentence describe a visual experience or scene? Visual
15. Does the input involve planning or organizing? Misc
16. Does the sentence involve spatial reasoning? Visual
17. Does the sentence involve an expression of personal values or beliefs? Abstract / Opinion
18. Does the sentence contain a negation? Abstract / Opinion
19. Does the sentence describe a sensory experience? Visual
20. Does the sentence include technical or specialized terminology? Misc
21. Does the input contain a number? Numeric
22. Does the sentence contain a cultural reference? Misc
23. Does the text describe a mode of communication? Social
24. Does the input include a comparison or metaphor? Abstract / Opinion
25. Does the sentence express a sense of belonging or connection to a place or community? Misc
26. Does the sentence describe a specific sensation or feeling? Misc
27. Does the text include a planning or decision-making process? Misc
28. Does the sentence include a personal anecdote or story? Misc
29. Does the sentence involve a discussion about personal or social values? Abstract / Opinion
30. Does the text describe a journey? Misc
31. Does the input contain a measurement? Numeric
32. Does the sentence describe a physical sensation? Misc
33. Does the sentence include a direct speech quotation? Social
34. Is the sentence reflective, involving self-analysis or introspection? Abstract / Opinion
35. Does the input describe a specific texture or sensation? Misc
36. Is the input about a discovery or realization? N/A
37. Does the sentence include an account of a miscommunication or misunderstanding? N/A
38. Does the sentence include a specific sound or auditory description? N/A
39. Does the sentence use a unique or unusual word? N/A
40. Does the sentence describe a change in a physical or emotional state? N/A
41. Does the sentence describe a moment of relief or resolution of tension? N/A
42. Does the sentence include a conditional clause? N/A
43. Does the sentence reference a specific time or date? N/A
44. Is the sentence conveying the narrator’s physical movement or action in detail? N/A
45. Is there mention of a city, country, or geographic feature? N/A
46. Does the sentence involve an unexpected incident or accident? N/A
47. Does the sentence involve a recount of a social or community event? N/A
48. Does the sentence express a philosophical or existential query or observation? N/A
49. Does the story involve a personal project or creation? N/A
50. Is the sentence emotionally positive? N/A
51. Does the sentence describe an activity related to daily life or routine? N/A
52. Does the text include a reference to a past era or time period? N/A
53. Does the input discuss a societal issue or social justice topic? N/A
54. Does the sentence convey a decision or choice made by the narrator? N/A
55. Does the sentence convey a sense of urgency or haste? N/A
56. Is the sentence providing an explanation or rationale? N/A